Tyrone is twice as old as his daughter. In six years Tyrone's age will be three times what his daughter's age was six years ago. How old are they at present? (We have been using the sublimation method and the addition method.)
Answer by scott8148(6628)

"Tyrone is twice as old as his daughter" ___ T = 2d

"In six years Tyrone's age will be three times what his daughter's age was six years ago" ___ T + 6 = 3(d - 6)

substituting (for the value of T) ___ 2d + 6 = 3d - 18 ___ 24 = d

substituting ___ T = 2(24) = 48